OpenStudy (anonymous):

Solve for x X/16 equals 7/10

OpenStudy (anonymous):

multiply both sides by 16 \[\color{red}{16}*\frac{ x }{ 16 } = \color{red}{16}*\frac{ 7 }{ 10 }\]

OpenStudy (anonymous):

that equals \[x = \frac{ 112 }{ 10 }\] divide them \[x = 11.2\]

OpenStudy (anonymous):

 set up the proportion and cross multiply, then solve for x